OpenStudy (anonymous):

What is the equation of y = x^3 with the given transformations? vertical stretch by a factor of 3, horizontal shift 4 units to the right, vertical shift 3 units down

OpenStudy (binary3i):

Y=3(x-4)^3 -3

OpenStudy (anonymous):

let f(x) = x^3 so with those transformations, your graph will be: y = 3f(x-4) + 3

OpenStudy (anonymous):

Hmm which one of you is right? xD

OpenStudy (binary3i):

mine

OpenStudy (binary3i):

see vertex will be at x=4 and that is suppose to be -3

OpenStudy (anonymous):

Okay thanks so much!

OpenStudy (anonymous):

were both right.... my answer is in terms of the transforms given... the other answer is the simplified form....

OpenStudy (binary3i):

when ever shifting is done we wright (y-k)=f(X-h)